OpenStudy (anonymous):

Perform the indicated operation g(x) = x3 - 2 h(x) = 2x - 1

OpenStudy (anonymous):

Perform the indicated operation g(x) = x3 - 2 h(x) = 2x - 1

OpenStudy (yamyam70):

are we solving for x here and substitute?

OpenStudy (anonymous):

(x)

OpenStudy (amoodarya):

g'=3x^2 g''=3*2x g''=3*(h(x)+1) ?